OpenStudy (giasimone):

Your friend gives you a gift card for $25.00. CDs cost $10.98 each plus 5.25% sales tax. You buy as many CDs as possible without having to pay any extra money. What is the balance on the card after the purchase? $

OpenStudy (texaschic101):

I am not sure about the correct way to do this problem....but I can tell you, if 2 CD's are bought....2 * 10.98 = 21.96 and 5.25% tax added is (.0525 * 21.96) = 1.16 making a total of 21.96 + 1.16 = 23.12. 25 - 23.12 = 1.88 left on the card

OpenStudy (giasimone):

alright thank you

OpenStudy (texaschic101):

sure thing :)

OpenStudy (giasimone):

could you help me on this one too

OpenStudy (giasimone):

You go to the wholesale club and buy a large bag of 12 smaller chip bags at $10.50. At your local store, the individual chip bags cost $0.99/bag. How much do you save per dozen bags by buying in bulk? $

OpenStudy (texaschic101):

at wholesale club, she got 1 dozen chip bags for $10.50 at local store, if she bought 1 doz. (12 bags) at .99 per bag, she would spend (12 x .99) = 11.88 11.88 - 10.50 = $1.38

OpenStudy (texaschic101):

you save $1.38 per dozen

OpenStudy (giasimone):

thank you